OpenStudy (anonymous):

Find x if the line through (-2, 3) and (x, 8) has a slope of 5/12

OpenStudy (anonymous):

Hi there, do you need help?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Do you know how to find the slope between 2 points?

OpenStudy (anonymous):

(8-3)/(x+2)=5/12. If you solve this equation, you will get the missing variable.

OpenStudy (anonymous):

Hmm. ok, and yes to the slope question

OpenStudy (anonymous):

Do you know how to solve the equation that I just posted?

OpenStudy (anonymous):

im working it out now

OpenStudy (anonymous):

I got x=4

OpenStudy (anonymous):



OpenStudy (anonymous):

That should help.

OpenStudy (anonymous):

yep.. thats right.. thank you!

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

can you help me with something else?

OpenStudy (anonymous):

Depends.

OpenStudy (anonymous):

ok, if you can, can you help with .. Write the equation of the line that has the indicated slope and contains the indicated point. Express the final equation in standard form. m = 3, (−5, 8)

OpenStudy (anonymous):

Use the point slope formula.

OpenStudy (anonymous):

I did, the trouble is converting it to standard form..

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

y=3x+23 ?

OpenStudy (anonymous):

That's y intercept form.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

y-8=3(x+5). That is after you enter the values into the point slope form.

OpenStudy (anonymous):

-3x+y=23

OpenStudy (anonymous):

I put that but its says its wrong. I got -3x+y=23 before

OpenStudy (anonymous):

Then rewrite it. 3x-y=-23

OpenStudy (anonymous):

it worked!

OpenStudy (anonymous):

i had the same problem with a question just like this

OpenStudy (anonymous):

for.. Write the equation of the line that contains the indicated pair of points. Express the final equation in standard form... I got -2x+y=3 but it says its wrong

OpenStudy (anonymous):

the points were (1, -1) (5, 7)

OpenStudy (anonymous):

Did you solve it correctly?

OpenStudy (anonymous):

I'm sure I did...

OpenStudy (anonymous):

What slope did you get?

OpenStudy (anonymous):

m=2

OpenStudy (anonymous):

okay, but your equation -2x+y=3 is not correct.

OpenStudy (anonymous):

it should be a 4, not a 3

OpenStudy (anonymous):

hmm.. ok so if I substitute the 4 everything is correct then?

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Actually, if it doesn't accept it, then rewrite it . 2x-y=-3

OpenStudy (anonymous):

I put that in before, and it said it was wrong

OpenStudy (anonymous):

i put it in before, 2x+y=-3, 2x-y=-3, y-2x=-3, and -2x+y=3 and they're all wrong... maybe I did the problem wrong

OpenStudy (anonymous):

Okay, what is the question asking? Just copy and paste it here.

OpenStudy (anonymous):

Write the equation of the line that contains the indicated pair of points. Express the final equation in standard form. (1, -1), (5, 7)

OpenStudy (anonymous):

ahh, This is the correct form. 2x-y=3

OpenStudy (anonymous):

Try that.

OpenStudy (anonymous):

Yes, it worked!

OpenStudy (anonymous):

I have 1 more that is just like this.. I need the correct standard form..

OpenStudy (anonymous):

Same question, but the new points are (-5, 4), (3, -3)

OpenStudy (anonymous):

Best response?

OpenStudy (anonymous):

well.. I got y=-7x-3/8

OpenStudy (anonymous):

thats not in standard of course, but is because its confusing

OpenStudy (anonymous):

just so you don't get confused